OpenStudy (anonymous):

Just need some help with these. I'll give a medal for the correct answer. 1. What is 9 1/4% of 50 2. What is 19% of $650 3. What is 3/8% of 4,000 4. 11 is 5% of what number? 5. 60 is 120% of what number 6. 43 is what percent of 78? 7. 22 is what percent of 80?

OpenStudy (triciaal):

percent means per 100 when working with percentages you are basically expressing everything as a fraction the whole is 100/100 1) 9 1/4% of 50 (9 1/4) / 100 =what number /50 9 1/4 =36/4 36/400 = x/50 solve for x the equivalent/50

OpenStudy (anonymous):

So 1 is 4.5?

OpenStudy (triciaal):

2) 19% of 650 19/100 x 650 =? 7) 22 is what percent of 80? x/100=22/80 what is x

OpenStudy (triciaal):

yes

OpenStudy (triciaal):

will check back

OpenStudy (anonymous):

I figured out most of them, but the last two are giving me trouble

OpenStudy (triciaal):

7) 22 is what percent of 80? x/100=22/80 what is x

OpenStudy (anonymous):

Thanks, I get it now

OpenStudy (triciaal):

great keep practicing welcome